# for Unseen Language Generalization

### v3

@@ Abstract
Pretrained language models (PLMs) have become remarkably adept at task and language generalization. Nonetheless, they often fail when faced with unseen languages. In this work, we present LinguAlchemy, a regularization method that incorporates various linguistic information covering typological, geographical, and phylogenetic features to align PLMs representation to the corresponding linguistic information on each language. Our LinguAlchemy significantly improves the performance of mBERT and XLM-R on low-resource languages in multiple downstream tasks such as intent classification, news classification, and semantic relatedness compared to fully finetuned models and displaying a high degree of unseen language generalization. We further introduce AlchemyScale and AlchemyTune, extension of LinguAlchemy which adjusts the linguistic regularization weights automatically, alleviating the need for hyperparameter search.
@@ 1 Introduction
Significant advancements in language processing technology have been achieved through the development of PLMs, leading to a commendable proficiency in language comprehension and generation Devlin et al. (2019); Liu et al. (2019); Lewis et al. (2019); Li et al. (2021); Sanh et al. (2022); Raffel et al. (2023). The advancement has been further expanded to non-English languages Conneau et al. (2020); Martin et al. (2020); Wilie et al. (2020); Kakwani et al. (2020); Cahyawijaya et al. (2021, 2024). However, there is still a gap in these models’ ability to generalize effectively to unseen languages. To mitigate this problem, there are some efforts to develop efficient language adaptation methods to unseen languages Pfeiffer et al. (2021b); Goyal et al. (2021); Alabi et al. (2022); Ebrahimi et al. (2022); Yong et al. (2023).
@@
Our strategy aims to refine cross-lingual generalization by leveraging linguistic features encapsulated in URIEL vectors. We hypothesize that languages with similar syntactic and typological characteristics can benefit from shared representational frameworks, significantly boosting performance in multilingual settings. This approach is particularly beneficial in contexts where language resources are limited.
@@
    LinguAlchemy is a new regularization approach that does not require any architectural change and can be adapted to different tasks and models.
@@ 2 Related Work
PLMs with their transformer-based architectures have been demonstrating exceptional capabilities in language comprehension and generation. Ganesh et al. (2021) have explored how these models learn intricate linguistic features, including syntax and semantics to enhance their performance across a wide range of language tasks Rathore et al. (2023).
@@
Incorporating new unseen languages has been a longstanding problem in the multilingual research, MAD-X Pfeiffer et al. (2020b) employ a language adapter to learn new unseen languages using language adapters that mitigate the risk of forgetting pre-trained knowledge, which is known as the curse-of-multilinguality Conneau et al. (2020). Nonetheless, this approach requires training for generalizing to new unseen languages, which makes it costly and difficult to scale to thousands of languages. MAD-G Ansell et al. (2021) and Udapter Üstün et al. (2020) further generalize this approach by utilizing a linguistic-driven contextual parameter generator (CPG) module to generate language-specific parameters, allowing the models to generalize to other languages with similar linguistic characteristics. Recently, Rathore et al. (2023) introduce ZGUL, which combines representations over multiple language adapters to generate the unseen language representation. Despite the effectiveness, all these approaches rely on two assumptions: (1) strict categorization of languages and (2) knowing the language category of the query apriori—our definition of “a priori categorization" as incorporating language-specific information into the model.
@@
In parallel, the development of linguistically-driven resources such as the URIEL vector and the lang2vec utility Littell et al. (2017) has been notable in extending multilingual NLP research, particularly for less-resourced languages. Complementing this, Ponti et al. (2019) emphasized the underutilization of typological features in existing databases and the need for integrating data-driven approaches of typological knowledge into language models. Previous studies have focused on extending multilingual NLP using URIEL vectors Lauscher et al. (2020); Lin et al. (2019); Tan et al. (2019); Oncevay et al. (2020), but none of their approaches use URIEL vectors for alignment during the finetuning process of language models.
@@
However, PLMs still face significant challenges in generalizing to unseen languages, particularly when adapting to low-resource and unseen languages. These challenges stem from the vast structural and semantic variation across languages Bender (2011); Jurafsky and Martin (2019), the scarcity of resources Mohammad (2019); Lewis et al. (2020), and the limitations inherent in the models themselves Lin et al. (2017). This situation highlights the complexity of scaling and generalizing these models effectively and underscores the need for more sophisticated approaches in model training and adaptation to ensure broader and more equitable language coverage.
@@ 3.2 Proof of Concept
2 Model Feature Type Acc.(%) mBERT geo 66.14 syntax_avg+geo 65.62 syntax_avg 66.41 syntax_knn+geo 66.08 syntax_knn 66.41 syntax_knn+syntax_avg+geo 66.47 syntax_knn+syntax_avg 66.41 XLM-R geo 80.16 syntax_avg+geo 80.54 syntax_avg 80.76 syntax_knn+geo 80.48 syntax_knn 80.32 syntax_knn+syntax_avg+geo 80.80 syntax_knn+syntax_avg 80.78
@@ 5.2 Effect of Scaling URIEL loss
Constant 10x 64.68% 80.32%
AlchemyScale 62.97% 80.43%
AlchemyTune 63.24% 79.10%
@@ 5.4 Seen Language Performance
While LinguAlchemy consistently improves performance across unseen languages, we note some inconsistencies concerning the performance of seen languages. In MASSIVE, we observe a noticeable performance drop in seen languages, while in contrast, we still see a massive gain in MasakhaNews and the performance of SemRel seems to be unaffected. The compiled results can be seen in Table 5.1.

### v4

@@ Abstract
Pretrained language models (PLMs) have become remarkably adept at task and language generalization. Nonetheless, they often fail when faced with unseen languages. In this work, we present LinguAlchemy, a regularization method that incorporates various linguistic information covering typological, geographical, and phylogenetic features to align PLMs representation to the corresponding linguistic information on each language. LinguAlchemy significantly improves the performance of mBERT and XLM-R on low-resource languages in multiple downstream tasks such as intent classification, news classification, and semantic relatedness compared to standard approach and displaying a high degree of unseen language generalization. We further introduce AlchemyScale and AlchemyTune, extension of LinguAlchemy which adjusts the linguistic regularization weights automatically, alleviating the need for hyperparameter search.
@@ 1 Introduction
Significant advancements in language processing technology have been achieved through the development of PLMs with their impressive capability in language comprehension and generation Devlin et al. (2019); Liu et al. (2019); Lewis et al. (2019); Li et al. (2021); Sanh et al. (2022); Raffel et al. (2023). The development has been further expanded to non-English languages Conneau et al. (2020); Martin et al. (2020); Wilie et al. (2020); Kakwani et al. (2020); Cahyawijaya et al. (2024). However, there is still a gap in these models’ ability to generalize effectively to low-resource and unseen languages, although there have been a numerous work in the field Pfeiffer et al. (2021b); Goyal et al. (2021); Alabi et al. (2022); Ebrahimi et al. (2022); Yong et al. (2023).
@@
Our strategy aims to refine cross-lingual generalization by leveraging linguistic features encapsulated in URIEL vectors. We hypothesize that languages with similar syntactic and geographical characteristics can benefit from shared representational frameworks, significantly boosting performance in multilingual settings. This approach is particularly beneficial in contexts where language resources are limited.
@@
    LinguAlchemy does not require any architectural change and can be adapted to different tasks and models.
@@ 2 Related Work
PLMs with their transformer-based architectures have been demonstrating exceptional capabilities in language comprehension and generation Ganesh et al. (2021). Rathore et al. (2023) have explored how these models learn intricate linguistic features, including syntax and semantics to enhance their performance across a wide range of language tasks.
@@
Incorporating new unseen languages has been a longstanding problem in the multilingual research, MAD-X Pfeiffer et al. (2020b) employ a language adapter to learn new unseen languages using language adapters that mitigate the risk of forgetting pre-trained knowledge, which is known as the curse-of-multilinguality Conneau et al. (2020). Nonetheless, this approach requires training for generalizing to new unseen languages, which makes it costly and difficult to scale to thousands of languages. MAD-G Ansell et al. (2021) and Udapter Üstün et al. (2020) further generalize this approach by utilizing a linguistic-driven contextual parameter generator (CPG) module to generate language-specific parameters, allowing the models to generalize to other languages with similar linguistic characteristics. Recently, Rathore et al. (2023) introduce ZGUL, which combines representations over multiple language adapters and adding linguistic vector information to generate the unseen language representation. Despite the effectiveness, all these approaches rely on two assumptions: (1) strict categorization of languages and (2) knowing the language category of the query apriori—our definition of “a priori categorization" as incorporating language-specific information into the model.
@@
In parallel, the development of linguistically-driven resources such as the URIEL vector and the lang2vec utility Littell et al. (2017) has been notable in extending multilingual NLP research, particularly for less-resourced languages, by providing methods to represent and compare the structured lingustic features across different languages. Complementing this, Ponti et al. (2019) pointed out the underexplored typological features in existing approaches and the need for integrating data-driven methods of typological knowledge into language models. Previous studies have focused on extending multilingual NLP using URIEL vectors Lauscher et al. (2020); Lin et al. (2019); Tan et al. (2019); Oncevay et al. (2020), but none of their approaches use URIEL vectors for alignment during the finetuning process of language models.
@@
However, PLMs still face significant challenges in generalizing to unseen languages, particularly when adapting to low-resource and unseen languages. These challenges stem from the vast structural and semantic variation across languages Bender (2011); Jurafsky and Martin (2019), the scarcity of resources Mohammad (2019); Lewis et al. (2020), and the limitations inherent in the models themselves Lin et al. (2017). This situation highlights the complexity of generalizing these models effectively to a broader and scope of language. This situation highlights the complexity of generalizing these models effectively to a broader scope of languages, making it difficult for them to perform well on languages they haven’t been specifically trained on.
@@ 3.2 Proof of Concept
Model Feature Type Acc.(%)
mBERT geo 66.14
syntax_avg+geo 65.62
syntax_avg 66.41
syntax_knn+geo 66.08
syntax_knn 66.41
syntax_knn+syntax_avg+geo 66.47
syntax_knn+syntax_avg 66.41
XLM-R geo 80.16
syntax_avg+geo 80.54
syntax_avg 80.76
syntax_knn+geo 80.48
syntax_knn 80.32
syntax_knn+syntax_avg+geo 80.80
syntax_knn+syntax_avg 80.78
@@ 5.2 Effect of Scaling URIEL loss
Constant 10x 64.68 80.32
AlchemyScale 62.97 80.43
AlchemyTune 63.24 79.10
@@ 5.4 Seen Language Performance
While LinguAlchemy consistently improves performance across unseen languages, we note some inconsistencies concerning the performance of seen languages. In MASSIVE, we observe a noticeable performance drop in seen languages, while in contrast, we still see a massive gain in MasakhaNews and the performance of SemRel seems to be unaffected. The compiled results can be seen in Table 4.